OpenStudy (anonymous):

Anna and Belle are asked to solve –2x – 15 = 6x + 9. Identify where one of them made an error. Anna Belle –2x – 15 = 6x + 9 –2x – 15 –6x = 6x + 9 –6x –8x – 15 = 9 –8x – 15 + 15 = 9 + 15 –8x = 24 Negative 8x over negative 8 equals 24 over negative 8 x = –3 –2x – 15 = 6x + 9 –2x – 15 + 2x = 6x + 9 + 2x –15 = 8x + 9 –15 + 9 = 8x + 9 + 9 –6 = 8x Negative 6x over 8 equals 8x over 8 Negative two-thirds

OpenStudy (anonymous):

1 Add 2x to both sides −15=6x+9+2x 2 Simplify 6x+9+2x to 8x+9 −15=8x+9 3 Subtract 9 from both sides −15−9=8x 4 Simplify −15−9 to −24 −24=8x 5 Divide both sides by 8 −248=x 6 Simplify 248 to 3 −3=x 7 Switch sides x=−3